### v1

# Group-in-Group Policy Optimization for LLM Agent Training

Abstract
Recent advances in group-based reinforcement learning (RL) have driven frontier large language models (LLMs) in single-turn tasks like mathematical reasoning. However, their scalability to long-horizon LLM agent training remains limited. Unlike static tasks, agent-environment interactions unfold over many steps and often yield sparse or delayed rewards, making credit assignment across individual steps significantly more challenging. In this work, we propose Group-in-Group Policy Optimization (GiGPO), a novel RL algorithm that achieves fine-grained credit assignment for LLM agents while preserving the appealing properties of group-based RL: critic-free, low memory, and stable convergence. GiGPO introduces a two-level structure for estimating relative advantage: (i) At the episode-level, GiGPO computes macro relative advantages based on groups of complete trajectories; (ii) At the step-level, GiGPO introduces an anchor state grouping mechanism that retroactively constructs step-level groups by identifying repeated environment states across trajectories. Actions stemming from the same state are grouped together, enabling micro relative advantage estimation. This hierarchical structure effectively captures both global trajectory quality and local step effectiveness without relying on auxiliary models or additional rollouts. We evaluate GiGPO on two challenging agent benchmarks, ALFWorld and WebShop, using Qwen2.5-1.5B-Instruct and Qwen2.5-7B-Instruct. Crucially, GiGPO delivers fine-grained per-step credit signals and achieves performance gains of > 12% on ALFWorld and > 9% on WebShop over the GRPO baseline: all while maintaining the same GPU memory overhead, identical LLM rollout, and incurring little to no additional time cost.††footnotetext: Code: https://github.com/langfengQ/verl-agent
1 Introduction
Large Language Models (LLMs) [1, 2, 3, 4] have leapt from static question-answer systems to versatile agents that perceive, reason, and act in open-ended environments. For instance, they now power embodied assistants that navigate simulated homes [5, 6], mobile and web navigators that plan multi-step browsing sessions [7, 8, 9], and autonomous explorers in interactive games [10, 11]. In these settings, LLM agents need to perceive, reason, and act in multi-turn loops, which requires not only language understanding but also long-horizon planning and decision-making.
Reinforcement learning (RL) [12] has become a crucial recipe for post-training LLMs, leading to frontier models like OpenAI o1 [13] and DeepSeek R1 [14]. In particular, group-based RL algorithms such as RLOO [15, 16] and GRPO [17] have proven especially effective in large-scale training. These methods replace value-function estimation with simple yet powerful relative advantage estimation within groups of rollouts. This group-based advantage computation enjoys favorable properties such as low memory overhead, critic-free optimization, and scalability to large models. However, their successes have so far been largely limited to single-turn tasks such as math problem solving [18, 19] and code generation [20], where reward arrives immediately and credit assignment is straightforward.
In contrast, LLM agents operating in external environments face fundamentally different learning landscapes. Their behavior unfolds over long episodes with tens of decision steps and tens of thousands of tokens (e.g., an ALFWorld [5] episode may include up to 50 steps and over 20k tokens). Rewards are typically sparse (sometimes arriving only at the end of an episode), and the impact of any individual action may only manifest much later in the trajectory. This substantially complicates the credit assignment for individual steps and increases the challenge of policy optimization. Naively applying existing group-based RL algorithms in such settings collapses step-level distinctions, undermining their effectiveness. Hence, these limitations raise a core question:
To address this, we introduce Group-in-Group Policy Optimization (GiGPO), a new group-based RL algorithm that nests two-dimensional notions of credit assignment, better-suited for the multi-step optimization of LLM agents. (i) At the episode level, GiGPO samples a group of complete trajectories under identical task and initial-state conditions, and computes macro relative advantages based on total returns like vanilla GRPO [17]. This captures the overall effectiveness of each trajectory and reflects the completeness of task execution. (ii) At the step level, GiGPO introduces a novel anchor state grouping mechanism for fine-grained relative advantages estimation. Specifically, it retroactively identifies the repeated environment states, termed as anchor states, across trajectories and uses them as anchors to construct step-level groups that allow for localized credit assignment.
The key insight behind GiGPO is that, under identical tasks and initial environment conditions, many trajectories within the group encounter the same states multiple times due to ineffective actions or loops, such as revisiting the same webpage, room, or game scene. These shared states provide a natural basis for step-level group construction and computing more granular advantage estimates. GiGPO uses these step-level groups to assign localized credit to actions based on their relative performance at a common state, enabling more precise optimization while avoiding the cost explosion of per-step extra rollouts. As such, GiGPO remains fully critic-free and requires no auxiliary value models while dramatically introducing finer credit signals for training LLM agents.
We evaluate GiGPO on two long-horizon agent benchmarks: ALFWorld [5], which tests embodied task planning in simulated household environments, and WebShop [21], which simulates complex, goal-driven web interactions. Our experiments with Qwen2.5-1.5B-Instruct and Qwen2.5-7B-Instruct [3] show that GiGPO consistently outperforms prompt-based agents, actor-critic baselines, and prior group-based RL methods. In particular, GiGPO injects fine-grained, step-level credit signals that sharpen policy learning of agents over long horizons and achieves performance improvement of > 12% on ALFWorld and > 9% on WebShop over GRPO. These gains come without compromising the core strengths of group-based RL (only < 0.002% time cost), making GiGPO a versatile and high-utility algorithm for the long-horizon optimization of LLM agents.
2 Related Work
LLMs as decision-making agents. The use of large language models (LLMs) as autonomous agents has expanded rapidly across domains such as program generation [22], smart device operation [23, 24, 25, 26], interactive gameplay [10], and robot behavior control [27]. Early works typically relied on leveraging pre-trained, frozen models through carefully designed prompting methods (like ReAct [28] and Reflexion [29]), enhanced memory and retrieval systems [11, 30], and integration with external tools [31, 32, 33]. More recent research has shifted toward adapting model parameters with supervised fine-tuning (SFT) [23] or RL [12], enabling agents to learn directly from environment interaction rather than static prompts or handcrafted workflows, which we introduce below.
Reinforcement learning for LLM agents. RL has played a pivotal role in enabling LLM agents to operate in dynamic, open-ended environments. Early work applied classical RL algorithms such as DQN [34] to train LLM agents in text-based games [35] and later research [36, 37, 38, 39, 40] started to employ value-based methods, such PPO [41] and AWR [42], in more diverse and interactive agent scenarios including Android device control [43], embodied ALFWorld [5], and card games [44]. More recent approaches have extended RL training to complex web-based and application-centered tasks. For instance, ArCHer [45] and AgentQ [46] target the WebShop benchmark [21], but require intricate designs and computation overhead such as additional value networks or Monte Carlo Tree Search (MCTS) [47]. Going further, LOOP [48] introduces a hybrid method combining REINFORCE leave-one-out (RLOO) [15, 16] with PPO-style updates, achieving state-of-the-art results in AppWorld [49]. RAGEN [50] introduces a trajectory-level GRPO that concatenates all states, intermediate reasoning, and actions into a unified episode-level response. However, it faces scalability challenges in long-horizon tasks (e.g., in ALFWorld, which involves up to 50 steps).
Reinforcement learning for large language models. An early and influential application of RL in LLMs is the Reinforcement Learning from Human Feedback (RLHF) [51, 52, 53, 54], which focuses on aligning LLMs to human preferences. Most recent works have explored using RL to enhance the capabilities of reasoning and logical deduction in LLMs [55, 14]. In particular, group-based RL algorithms have emerged as promising alternatives to traditional methods like PPO [41]. These methods, such as RLOO [15, 16], GRPO [17], Dr. GRPO [18], DAPO [19], and CPPO [56], avoid introducing extra value functions by leveraging a group of samples from the same query and estimate the advantages accordingly. This enables the large-scale RL training and has shown strong results in tasks such as mathematical reasoning [14], search [57, 58], and tool use [59, 60]. Our work is closely related to this line of research, with a focus on training LLM agents. We aim to retain the benefits of group-based RL, such as critic-free learning and efficiency, while introducing finer-grained credit assignment. Moreover, the hierarchical core of GiGPO is orthogonal to existing group-based RL approaches, making it fully compatible and capable of incorporating them to enhance performance.
3 Preliminaries
Problem setup. We consider a general setting in which an LLM agent interacts with an environment to complete multi-step tasks based on a task description . At each discrete time step , the agent observes a state and generates a textual action , where denotes the token vocabulary and is the maximum generation length. The environment then returns a scalar reward and the next state . A full episode consists of a trajectory . The agent’s behavior is governed by an LLM policy , parameterized by , which defines a distribution over outputs conditioned on the current state and the task prompt . In many realistic scenarios, the environment may provide sparse or delayed rewards (e.g., success and failure signals at the end of an episode) or weak feedback signals for intermediate steps. As the agent generates consecutive textual actions (), each potentially spanning thousands of tokens, it becomes particularly challenging to assign credit to individual tokens over the course of an episode.
Group-based RL. Recent RL works converge on a simple recipe for training LLMs: for a given task description , the LLM samples a group of candidate trajectories , each corresponding to one full episode rollout under . Each trajectory receives a scalar reward reflecting the overall quality or success of the generated outcome. Instead of estimating advantages using separate value functions like PPO [41], group-based RL methods compute advantages purely based on the statistics internal to the sampled group:
For example, in GRPO [17], the advantage of each trajectory is estimated by normalizing its reward with respect to the group’s mean and standard deviation. This design is highly memory-efficient and can scale effectively to the large batch sizes and model sizes typical in modern LLM training, making it a practical and scalable choice for large-scale RL training.
4 Training LLM Agents with GiGPO
While group-based RL algorithms [17, 14] have proven highly effective for training LLMs in single-turn tasks, their extension to multi-step agent settings faces critical challenges in credit assignment. Figure 1 illustrates this gap. Vanilla GRPO (left) treats each trajectory as a whole and computes a single relative advantage for the entire episode, which fails to provide actionable feedback for individual steps. A natural remedy is to roll out multiple single-step actions for each state via as shown in Figure 1 (middle). However, this approach quickly becomes impractical due to the substantial overhead of extra LLM forward passes and the difficulty of evaluating rewards for hypothetical actions never actually executed.
To overcome these challenges, we propose our Group-in-Group Policy Optimization (GiGPO) in this section. Similar to prior works [48, 50], GiGPO begins by sampling groups of trajectories under identical tasks and initial environment states. It then introduces a two-level grouping structure: preserving episode-level grouping for holistic performance comparison, while dynamically constructing an additional set of step-level groups by retroactively aggregating actions encountering the same environment states. This “group-in-group” construction yields two complementary advantages: (1) Episode relative advantages capture the holistic effectiveness of each trajectory, providing a stable, global training signal. (2) Step relative advantages zoom in on which actions outperform their peers within the same state, endowing the gradient with fine-grained credit.
Figure 2 presents an overview of the GiGPO training pipeline. In the remainder of this section, we will detail the computation of episode-level relative advantages, elaborate on the anchor state grouping mechanism, describe the derivation of step-level relative advantages, and finally present the overall GiGPO objective.
4.1 Episode Relative Advantages
We first introduce the episode-level relative advantages, which represent the coarse-grained component of GiGPO and mirror the naive application of GRPO at the trajectory level. We roll out the agent’s policy in the environment to collect complete trajectories under a fixed task and identical initial states. Formally, this process yields a group of trajectories , where each trajectory is denoted as and the initial states satisfy . For each trajectory, we utilize the total return as a holistic measure of how effectively the agent completes the task. In settings where only a binary reward is given at the end of the episode, this simplifies to for success and for failure. Then, we organize the trajectories and their corresponding returns into an episode-level group:
To evaluate the global relative quality of each trajectory within the group, we compute an episode relative advantage for each by normalizing the total return with the group’s mean and a normalization factor:
In GRPO [17], the default normalization factor is defined as the standard deviation, i.e., . However, this may introduce a difficulty bias [18], where trajectories from low-variance groups (e.g., very easy or hard tasks) receive disproportionately large gradients. In the context of the LLM agent, where tasks often involve very long horizons, this effect tends to emerge frequently, potentially affecting the stability of updates. As an alternative, we also consider a fixed normalization factor , which yields an unbiased Leave-One-Out estimator [15] (see Appendix B for details). This simple adjustment helps stabilize training in some challenging agent scenarios.
Overall, the episode relative advantage captures whether the agent successfully completes the assignment across the entire decision horizon . Similar to the vanilla GRPO for multi-step optimization shown in Figure 1 (left), it primarily focuses on macro credit assignment, encouraging the policy to develop coherent, trajectory-wide behaviors that maximize overall task performance.
4.2 Step Relative Advantages
While the episode relative advantage offers a macro, trajectory-wide signal, it cannot distinguish between the contributions of individual actions within the trajectory. To obtain this fine-grained feedback, we need to form step-level groups: for the same state, we gather the different actions and compare their outcomes, thereby learning which choices are relatively better or worse. A naive way to do so would be to roll out fresh actions from every state (Figure 1, middle), but that is prohibitively expensive. Instead, we introduce anchor state grouping below, avoiding extra LLM overhead.
Anchor state grouping. As all trajectories arise from the same task and identical initial conditions, many environment states naturally recur across episodes and even across time steps within a single trajectory. We leverage this redundancy by identifying and grouping identical states across trajectories, thereby dynamically constructing step-level groups. Specifically, let denote the set of all distinct environment states appearing in the trajectory group . We treat each such unique state as an implicit anchor and use it to gather all matching occurrences of that state, and therefore call as “anchor state”. Based on this, we can construct step-level groups (one for each unique anchor state ), which is defined as follows:
Unlike per-state rollout, this procedure incurs no extra rollouts: it is entirely offline and requires only lightweight key-based grouping using hashmaps. Each group contains multiple instances of the same environment state paired with potentially different actions. Hence, this structure effectively builds the step-level group, forming the basis for subsequent step-level advantage estimation.
Relative advantage computation. Although each tuple contains an immediate reward , it may be sparse, especially in long-horizon tasks. To better capture long-term impact, we associate a discounted return with each step. Let be the standard RL discount factor. For each element in , we compute its discounted return by
This quantity captures the future impact of action on subsequent rewards, rather than relying solely on the immediate reward . Accordingly, the step-level group for each becomes:
Once these step-level groups are formed, we compute the step relative advantage for each and each action in :
provides micro credit assignment and fine-grained feedback on the relative quality of individual actions taken from the same state. In contrast to the coarse, trajectory-wide signal of , it offers step-level guidance that is essential for refining decisions in long-horizon agent tasks.
How does step-level group work? We present an intuitive illustration in Figure 3 to show the utility of the step relative advantages. We consider two example trajectories from the set . In , the agent first selects the 2nd Item (incorrect), then returns to the previous page and selects the 1st Item (correct), successfully completing the task. Due to temporal discounting (Equation (5)), the earlier action (2nd Item) receives a lower discounted return than the later correct one (1st Item). In , the agent clicks the Next Page, ultimately failing to find the target and receiving no reward. By aggregating these actions into the same step-level group based on their shared anchor state, GiGPO computes their relative advantages and yields a clear preference ordering: . This ranking successfully captures fine-grained distinctions in long-term utility that are missed by prior group-based RL methods [16, 17, 19]. While this example illustrates GiGPO’s effectiveness in sparse-reward environments, its advantages extend naturally to dense-reward scenarios, where per-step rewards can be fully leveraged to assess the relative quality of individual actions within shared states.
4.3 Group-in-Group Policy Optimization
We finally combine the two levels of advantage signals into a single group-in-group advantage to assign credit at both global (episode) and local (step) scales:
where is a weighting coefficient that balances episode relative advantage and step relative advantage. captures how good an episode is compared to others in the group, while refines step-level performance within shared environment state conditions. Jointly, they provide hierarchical supervision for the policy optimization of LLM agents. Then the clipped policy optimization objective of GiGPO is:
where is the importance sampling ratio, controls the strength of the KL penalty encouraging proximity to a reference policy . We present the pseudo code in Appendix C.
5 Experiment
In this section, we present empirical evaluations of GiGPO across a variety of LLM agent tasks. Specifically, we aim to demonstrate: (1) the strong ability of GiGPO in training LLM agents; (2) the ablation study of GiGPO; (3) the dynamic trend of step-level group over the course of training; (4) the computational budget of GiGPO.
5.1 Experiment Setup
Environments. We train the LLM agents on two challenging benchmarks: ALFWorld [5] and WebShop [21]. ALFWorld is an embodied environment designed to assess the ability of LLM agents to perform multi-step decision-making. In each episode, the agent receives a text goal and must accomplish it through multi-turn interaction with the environment. It includes 4,639 task instances across six categories of common household activities: Pick & Place (Pick), Examine in Light (Look), Clean & Place (Clean), Heat & Place (Heat), Cool & Place (Cool), and Pick Two & Place (Pick2). WebShop is a complex, web-based interactive environment designed to test the LLM agents in realistic online shopping scenarios. To complete the task, the agent must interact with a simulated HTML-based shopping website to search for, navigate to, and ultimately purchase a suitable item. It contains over 1.1 million products and 12k user instructions, providing a rich and diverse action space.
Baselines. We compare our approach with a range of competitive baselines: (1) Closed-source LLMs: GPT-4o [1] and Gemini-2.5-Pro [2], which represent state-of-the-art capabilities in general-purpose reasoning and language understanding. (2) Prompting agents: ReAct [28] and Reflexion [29], which rely on in-context prompting to guide multi-step behavior without parameter updates. (3) RL training methods: PPO [41], a widely used actor-critic algorithm that requires an additional value model, and group-based critic-free methods RLOO [15, 16] and GRPO [17], which perform advantage estimation over trajectory groups.
Training details. We use Qwen2.5-1.5B-Instruct and Qwen2.5-7B-Instruct [3] as our base models. The weighting coefficient is set to with no further tuning. All RL training methods (including ours and the baselines) use exactly the same hyperparameter configurations. The rollout group size for group-based RL methods is set to 8. The LLM agent is prompted to first generate a chain-of-thought [61] enclosed within <think> </think> tags, followed by the action enclosed within <action> </action> tags. Full training settings and hyperparameter details are provided in Appendix D.1.
5.2 Performance
| Type | Method | ALFWorld | WebShop | |||||||
| Pick | Look | Clean | Heat | Cool | Pick2 | All | Score | Succ. | ||
| Base: Closed-Source Model | ||||||||||
| Prompting | GPT-4o | 75.3 | 60.8 | 31.2 | 56.7 | 21.6 | 49.8 | 48.0 | 31.8 | 23.7 |
| Prompting | Gemini-2.5-Pro | 92.8 | 63.3 | 62.1 | 69.0 | 26.6 | 58.7 | 60.3 | 42.5 | 35.9 |
| Base: Qwen2.5-1.5B-Instruct | ||||||||||
| Prompting | Qwen2.5 | 5.9 | 5.5 | 3.3 | 9.7 | 4.2 | 0.0 | 4.1 | 23.1 | 5.2 |
| Prompting | ReAct | 17.4 | 20.5 | 15.7 | 6.2 | 7.7 | 2.0 | 12.8 | 40.1 | 11.3 |
| Prompting | Reflexion | 35.3 | 22.2 | 21.7 | 13.6 | 19.4 | 3.7 | 21.8 | 55.8 | 21.9 |
| RL Training | PPO (with critic) | 64.8±3.5 | 40.5±6.9 | 57.1±4.9 | 60.6±6.6 | 46.4±4.0 | 47.4±1.9 | 54.4±3.1 | 73.8±3.0 | 51.5±2.9 |
| RL Training | RLOO | 88.3±3.0 | 52.8±8.6 | 71.0±5.9 | 62.8±8.7 | 66.4±5.5 | 56.9±4.7 | 69.7±2.5 | 73.9±5.6 | 52.1±6.7 |
| RL Training | GRPO | 85.3±1.5 | 53.7±8.0 | 84.5±6.8 | 78.2±7.9 | 59.7±5.0 | 53.5±5.6 | 72.8±3.6 | 75.8±3.5 | 56.8±3.8 |
| RL Training | GiGPOw/ std | 94.4±5.9 | 67.5±4.6 | 94.8±3.8 | 94.4±7.8 | 79.8±4.7 | 76.4±5.4 | 86.7±1.7 | 83.1±1.6 | 65.0±3.2 |
| RL Training | GiGPOw/o std | 96.0±1.4 | 76.5±3.9 | 91.8±5.5 | 91.3±6.3 | 71.7±8.4 | 79.5±7.7 | 86.1±4.7 | 83.5±1.8 | 67.4±4.5 |
| Base: Qwen2.5-7B-Instruct | ||||||||||
| Prompting | Qwen2.5 | 33.4 | 21.6 | 19.3 | 6.9 | 2.8 | 3.2 | 14.8 | 26.4 | 7.8 |
| Prompting | ReAct | 48.5 | 35.4 | 34.3 | 13.2 | 18.2 | 17.6 | 31.2 | 46.2 | 19.5 |
| Prompting | Reflexion | 62.0 | 41.6 | 44.9 | 30.9 | 36.3 | 23.8 | 42.7 | 58.1 | 28.8 |
| RL Training | PPO (with critic) | 92.3±4.0 | 64.0±8.4 | 92.5±2.4 | 89.5±7.0 | 80.3±2.0 | 68.8±8.3 | 80.4±2.7 | 81.4±3.1 | 68.7±5.1 |
| RL Training | RLOO | 87.6±4.3 | 78.2±8.3 | 87.3±5.8 | 81.3±7.6 | 71.9±5.2 | 48.9±8.4 | 75.5±4.6 | 80.3±3.2 | 65.7±4.0 |
| RL Training | GRPO | 90.8±5.1 | 66.1±6.7 | 89.3±5.4 | 74.7±6.9 | 72.5±5.4 | 64.7±7.3 | 77.6±5.2 | 79.3±2.8 | 66.1±3.7 |
| RL Training | GiGPOw/ std | 97.7±1.6 | 82.7±7.9 | 98.8±1.6 | 83.7±7.2 | 89.3±8.2 | 79.2±6.6 | 90.8±1.3 | 84.4±2.9 | 72.8±3.2 |
| RL Training | GiGPOw/o std | 91.8±5.4 | 88.6±6.3 | 95.9±3.2 | 90.2±2.6 | 86.5±5.5 | 85.2±7.5 | 90.2±2.3 | 86.2±2.6 | 75.2±3.8 |
Table 1 demonstrates the strong performance of GiGPO across both ALFWorld and WebShop. As shown, Closed-source LLMs offer only moderate performance: Gemini-2.5-Pro reaches 60.3% success on ALFWorld and 35.9% on WebShop, while GPT-4o lags further behind. Open-source prompt-only agents (e.g., ReAct and Reflexion) show marginal improvements over vanilla prompting but still underperform, underscoring the difficulty of long-horizon control without post-training. RL training brings substantial gains: PPO improves average ALFWorld success to 54.4% on the 1.5B model and 80.4% on the 7B model, with WebShop scores also increasing significantly. However, this comes at the expense of increased complexity: requiring a separate critic network, hyperparameter tuning, and longer training durations [62, 48]. GRPO and RLOO also yield strong performance while being more computationally efficient, demonstrating the effectiveness of group-based RL in large-scale LLM training. Nevertheless, their lack of fine-grained per-step feedback limits their ability to provide precise credit assignment across long horizons. In contrast, GiGPO overcomes this limitation with a two-level advantage estimation, enabling both GiGPOw/ std and GiGPOw/o std to consistently surpass GRPO and RLOO. In particular, GiGPOw/o std surpasses GRPO by 13.3% on ALFWorld and 10.6% on WebShop at 1.5B, and by 12.6% and 9.1%, respectively, at 7B. These results highlight GiGPO’s superior ability to train LLM agents more effectively and more efficiently. We also find that GiGPO enables agents to exhibit emergent reasoning behavior (see Appendix E).
Lastly, we observe that the normalization factor is task-dependent rather than universally helpful. On relatively difficult tasks (such as Look, Pick2, and WebShop), standard-deviation scaling () could exaggerate gradients from overly difficult samples or highly imbalanced groups, harming update stability; fixing therefore yields higher success. Yet, offers no clear advantage on other tasks and both variants perform similarly, which aligns with findings in [63]. This suggests that can still be beneficial when reward variance is stable.
5.3 Ablation Study
Next, we conducted an ablation study, comparing GiGPOw/o std (), GiGPOw/ std (), GiGPOw/o (without step relative advantages), and GiGPOw/o (without episode relative advantages) to evaluate the impact of each component on performance. We use Qwen2.5-1.5B-Instruct as the policy of the agent. The results are presented in Figure 4.
As illustrated, eliminating either component of the two-level advantage significantly degrades performance. Removing the episode relative advantages (GiGPOw/o ) leads to a substantial drop across all tasks, as the policy no longer receives a stable, trajectory-wide signal to encourage long-range coherence. Similarly, discarding the step relative advantages (GiGPOw/o ) results in pronounced declines, particularly on more complex and demanding tasks such as Cool, Pick2, and WebShop, which may require nuanced training feedback at each decision step. In such cases, precise per-step credit assignment is essential for effective learning and policy refinement. Moreover, we can see that the relative performance gap between GiGPOw/ std and GiGPOw/o std is comparatively minor compared to that observed in structural ablations. This suggests that the combination of episode- and step-level signals is the primary driver of performance gains, and that each component is crucial for training LLM agents effectively.
5.4 Dynamics of Step-Level Group
In this part, we examine how the distribution of step-level groups evolves throughout training to better understand the utility of GiGPO. We use Qwen2.5-1.5B-Instruct as the base model. We train the LLM agent in ALFWorld and track changes in step-level group sizes throughout training.
As shown in Figure 5, we observe that step-level groups of size 1 (i.e., those with ) only account for < 35% throughout training. This indicates that the majority of states (over 65%) recur across trajectories and therefore contribute to achor state grouping. Moreover, at iteration 10, large group sizes account for over 20%, reflecting behavioral redundancy in the early stages of training. This is consistent with the fact that immature policies often produce invalid actions or fall into repetitive loops. As training progresses, the group size distribution changes markedly. By iteration 75, we observe a substantial reduction in extreme group sizes: drops from 16.2% to 12.1% and drops from 5.6% to 3.1%. This shift suggests that the agent is learning to avoid previously common dead ends and invalid actions, and is beginning to exhibit more diverse and purposeful decision-making. At iteration 140, the distribution becomes tightly concentrated around group sizes of 6 to 8. Given that is set to , this convergence implies that all 8 trajectories within an episode group are now behaving consistently: LLM agent has learned a coherent and robust policy for completing the task, aligning with the plateau in success rate ().
5.5 Computational Budget
Lastly, we analyze the computational budget of GiGPO. GiGPO shares the same core architecture as GRPO, including multi-turn rollouts, computation of old and reference probabilities, and clipped policy updates. Both approaches are critic-free and operate with a single actor LLM, thereby resulting in identical GPU memory usage and LLM rollout costs. The primary additions introduced by GiGPO are the step-relative advantage estimation components, as described in Section 4.2. To evaluate their costs, we train an LLM agent in ALFWorld using Qwen2.5-1.5B-Instruct and record a detailed breakdown of per-iteration training time.
As shown in Figure 6, the additional components incur little to no additional time cost compared to dominant operations such as rollouts, computation of old and reference probabilities, and policy updates, whose total time cost reaches 362.83s per iteration. In contrast, anchor state grouping (involving hashmap lookups) takes only 0.01s per iteration, and the step-relative advantage computation (involving simple arithmetic) adds just 0.53s. These operations account for < 0.002% of the total per-iteration training time, demonstrating that GiGPO shares the same high computational efficiency as GRPO.
6 Conclusions and Limitations
In this work, we proposed GiGPO, a novel group-based RL algorithm to tackle the credit assignment challenge in long-horizon LLM agent training. GiGPO introduces a hierarchical advantage estimation that enables fine-grained per-step credit assignment while retaining the efficiency and stability of group-based RL. By retroactively grouping steps that share the same state across trajectories, it achieves this without incurring additional LLM rollout or GPU memory overhead. Empirical evaluations across two complex environments, ALFWorld and WebShop, demonstrate that GiGPO significantly outperforms both prompt-based agents and prior RL methods. A potential limitation of GiGPO is its reliance on state matching for anchor group construction. In highly complex environments, identical states may be hard to detect due to noise or subtle differences. Despite this, GiGPO still retains a strong performance lower bound: in the extreme case where no states are repeated across trajectories (i.e., ), it naturally degrades to GRPO, preserving GRPO’s effectiveness and stability in credit assignment. However, we suggest a more robust solution: introducing state similarity through embeddings or approximate matching, which could better capture structurally equivalent states. We leave this exploration as a promising direction for future work.
Appendix A Source Code
As part of the new assets released with this work, we provide the complete source code at https://github.com/langfengQ/verl-agent. The repository includes a comprehensive README.md with installation instructions, training and evaluation procedures, and license information.
Our code is built upon the veRL framework [64] and extended with several features to support scalable training of long-horizon LLM agents. Key capabilities of our code include: (1) a step-wise multi-turn interaction paradigm that avoids concatenating the entire interaction history into a single input, as done in RAGEN [50], thereby ensuring scalability for very long-horizon optimization, (2) support for parallelized, gym-style group-based environments to enable high-throughput and state-aligned rollouts, (3) a rich suite of environments, including ALFWorld, WebShop, Sokoban, and Gym Cards; and (4) compatibility with both text-based and vision-language agents.
We provide a detailed installation guide (setup takes approximately five minutes), along with out-of-the-box scripts to reproduce our main results in ALFWorld and WebShop, as well as additional experiments on Sokoban and Gym Cards presented in Appendix D.3.
Appendix B Unbiasedness
We show that setting leads to an unbiased estimator up to a constant scaling factor. Our derivation follows the approach in [18]. For comparison, the standard REINFORCE Leave-One-Out (RLOO) [15, 16] is defined as:
We can relate () and by introducing a scaling factor of :
Thus, setting corresponds to a rescaled version of and scaling the advantage by a constant does not affect the dynamics of policy gradient (it can be absorbed into the learning rate).
Appendix C Pseudo Code
Algorithm 1 summarizes the full GiGPO training procedure. Compared to vanilla GRPO, we highlight the additional parts introduced by GiGPO in italics. In particular, building step-level groups is implemented by treating anchor states as keys and aggregating corresponding data into a hash table, which incurs minimal overhead. Furthermore, computing step relative advantages and combining advantages involve only simple arithmetic operations, both of which are lightweight. As such, GiGPO preserves the critic-free, low-memory, and stable convergence properties of group-based RL, while introducing fine-grained credit assignment that is particularly beneficial for training long-horizon LLM agents.
Appendix D Experiment Details
D.1 Details of Training
Hyperparameters for ALFWorld. All methods are configured with identical hyperparameters: the maximum prompt length is 2048 tokens, and the maximum response length is 512 tokens. Each episode allows up to 50 environment steps. The learning rate is set to 1e-6 for the actor and 1e-5 for the critic (used only in PPO). We adopt a rule-based reward, assigning a reward of 10 for success and 0 for failure. To handle invalid actions generated by the agent, we apply a reward penalty of -0.1. For all group-based RL methods, we use a group size of 8 and sample 16 different groups per rollout, resulting in a total of environments. In contrast, PPO uses 128 separate environments for rollouts. The rollout temperature is set to 1.0, while the validation temperature is set to 0.4. The mini-batch size is 256, and the KL-divergence loss coefficient is set to 0.01. For GiGPO, the weighting coefficient is fixed at 1 without further tuning, and the discount factor is set to 0.95.
Hyperparameters for WebShop. All methods are configured with identical hyperparameters: the maximum prompt length is 4096 tokens, and the maximum response length is 512 tokens. Each episode is limited to 15 environment steps. The learning rate is 1e-6 for the actor and 1e-5 for the critic (used only in PPO). We adopt a rule-based reward, assigning a reward of 10 for success and 0 for failure. Invalid actions are penalized with a reward of -0.1. As with ALFWorld, all group-based RL methods use a group size of 8 and sample 16 groups per rollout, totaling environments. PPO, on the other hand, uses 128 distinct environments for rollouts. The rollout temperature is set to 1.0, while the validation temperature is set to 0.4. The mini-batch size is 64, and the KL-divergence loss coefficient is set to 0.01. For GiGPO, the weighting coefficient is set to 1 without additional tuning, and the discount factor is set to 0.95.
Computing Details. Experiments using Qwen2.5-1.5B-Instruct are conducted on two NVIDIA H100 GPUs, while those using Qwen2.5-7B-Instruct are trained on four NVIDIA H100 GPUs. Each experiment is trained for a total of 150 training iterations.
D.2 Prompts
The prompts we use for LLM agents are presented in Figure 7 and Figure 8. These prompt templates are constructed using Python-style string formatting, where placeholders enclosed in curly braces ({}) represent semantic slots. These placeholders, such as {task_description}, {step_count}, and {current_observation}, are dynamically populated at runtime via Python’s .format() function. To enrich the agent’s context, we use historical information and set the history length to 2. A fully instantiated example is shown in Figure 9 and Figure 10. This augmented representation is used as the final input to the agent.
The <think>…</think> block instructs the agent to explicitly perform step-by-step reasoning, thereby promoting chain-of-thought style deliberation. The <action>…</action> block is used to clearly indicate the final action decision.
D.3 Performance on Vision-Language Agents
We conduct additional experiments in vision-language model (VLM) settings where agents must reason over both visual and textual inputs. We use two interactive game environments: Sokoban [65] with a 66 size and EZPoints in Gym Cards [38] as shown in Figure 11. Sokoban is a classic grid-based puzzle game where the agent must push boxes onto designated goal positions. Solving Sokoban requires spatial reasoning and long-term planning. In EZPoints, the agent is presented with two playing cards and must construct an arithmetic formula step-by-step such that the final result equals a target value of 12. This setting emphasizes symbolic reasoning and multi-step planning. In these tasks, the agent receives an RGB image representing the current environment state along with a textual instruction, and must take sequential actions to complete the task. All methods are built upon Qwen2.5-VL-3B-Instruct [66].
As shown in Table 2, GiGPO significantly outperforms the prompting baseline and GRPO on both tasks. Notably, GiGPO achieves 81.0% success on Sokoban and 100% success on EZPoints, suggesting its ability to generalize beyond language-only settings.
| Type | Method | Sokoban [66] | EZPoints |
| Prompting | Qwen2.5-VL | 11.7 | 3.1 |
| RL Training | GRPO | 67.1±4.7 | 86.9±3.4 |
| RL Training | GiGPOw/ std | 76.9±2.7 | 100.0±0.0 |
| RL Training | GiGPOw/o std | 81.0±3.6 | 100.0±0.0 |
D.4 Orthogonality to Single-Turn Group-Based RL
GiGPO remains orthogonal to other advancements in single-turn group-based RL, allowing it to incorporate complementary techniques without loss of generality. To verify this, we consider DAPO [19], which builds on GRPO with dynamic sampling and clip-higher techniques. We integrate both into GiGPO, yielding a variant denoted as GiGPOdynamic.
As shown in Table 3, DAPO improves over GRPO, confirming the benefits of its techniques. Importantly, GiGPOdynamic further outperforms DAPO, demonstrating that it can effectively benefit from and amplify such improvements, achieving better results (75.0%) on WebShop.
| Type | Method | Score | Success Rate (%) |
| Prompting | Qwen2.5 | 23.1 | 5.2 |
| RL Training | GRPO | 75.8±3.5 | 56.8±3.8 |
| RL Training | DAPO | 84.6±2.9 | 66.1±3.2 |
| RL Training | GiGPOdynamic | 87.5±1.6 | 75.0±3.5 |
Appendix E Reasoning Behavior
We next present the complete trajectory of multi-turn interaction between agent and environment, and demonstrate the reasoning behavior of LLM agents (Qwen2.5-7B-Instruct) trained by GiGPO.